The population of a town grows at a rate proportional to the population present at time t. The initial population of 5000 increases by 50% in 10 years. What will be the population in 30 years?


1
Expert's answer
2021-03-01T07:58:57-0500

According to the question,

"\frac{dP}{dt}=kP", where "P" is population at any time

solving equation,

"\int\frac{dP}{P}=\int kdt" ,

"lnP=kt+C,"

Applying the conditions,

at "t=0, \space P=5000"

"ln(5000)=C"

at "t=10,\space P=7500"

"ln(7500)=k(10)+ln(500)"

"k=\frac{ln(1.5)}{10}=0,040547"

So Equation of the Population growth is given by

"P=5000e^{0.040547t}"

Population in 30 years will be

"P=5000e^{0.040547 \cdot30}=16875.25\approx 16875" persons